OpenStudy (lisa123):

simplify (|pi-2^3|)/(8-pi)- 3+14÷2-5

OpenStudy (lisa123):

its a pemdas type of problem

OpenStudy (lisa123):

the answer is -4 I just don't know how to get it

OpenStudy (decentnabeel):

first slove the parenthesis then apply DMAS rule

OpenStudy (decentnabeel):

ok

OpenStudy (decentnabeel):

\[\frac{14}{2}-3-5+\frac{\left|\pi -2^3\right|}{-\pi +8}\] \[\mathrm{Divide\:the\:numbers:}\:\frac{14}{2}=7\] \[=\frac{\left|\pi -2^3\right|}{-\pi +8}-3-5+7\] \[\mathrm{Add/Subtract\:the\:numbers:}\:-3-5+7=-1\] \[=\frac{\left|\pi -2^3\right|}{-\pi +8}-\] \[\mathrm{Combine\:the\:fractions\:using\:the\:LCD}:\quad \frac{\left|\pi -2^3\right|}{-\pi +8}-1=\frac{-\left(-\pi +8\right)+\left|\pi -2^3\right|}{-\pi +8}\] \[=\frac{-\left(-\pi +8\right)+\left|\pi -2^3\right|}{-\pi +8}\] \[=\frac{-\left(-\pi +8\right)+\left|\pi -8\right|}{-\pi +8}\] \[\mathrm{Expand}\:-\left(-\pi +8\right):\quad \pi -8\] \[=\frac{\left|\pi -8\right|-8+\pi }{-\pi +8}\] \[\frac{14}{2}-3-5+\frac{\left|\pi -2^3\right|}{-\pi +8}=\frac{\left|\pi -8\right|-8+\pi }{-\pi +8}\quad \left(\mathrm{Decimal:\quad }\:0\right)\]

OpenStudy (decentnabeel):

ok @lisa123

OpenStudy (lisa123):

ok but why is the answer not -4?

OpenStudy (decentnabeel):

the answer is 0 not -4

OpenStudy (decentnabeel):

or you sure the answer is -4 @lisa123

OpenStudy (lisa123):

that's what I got on mathway

OpenStudy (lisa123):

oh Im soooo sorry guys! theres a bracket in -(3+14 div 2 -5). I have to distribute the negative right?

OpenStudy (decentnabeel):

\[-\left(\frac{14}{2}+3-5\right)+\frac{\left|\pi -2^3\right|}{-\pi +8}=\frac{-5\left(-\pi +8\right)+\left|\pi -8\right|}{-\pi +8}\quad \left(\mathrm{Decimal:\quad }\:-4\right)\]

OpenStudy (lisa123):

oh ok! thank you @DecentNabeel and @LynFran

OpenStudy (decentnabeel):

\[\mathrm{Simplify}\:-\left(\frac{14}{2}+3-5\right):\quad -5\] \[=\frac{\left|\pi -2^3\right|}{-\pi +8}-5\] \[\mathrm{Combine\:the\:fractions\:using\:the\:LCD}:\quad \frac{\left|\pi -2^3\right|}{-\pi +8}-5=\frac{-5\left(-\pi +8\right)+\left|\pi -2^3\right|}{-\pi +8}\] \[=\frac{-5\left(-\pi +8\right)+\left|\pi -2^3\right|}{-\pi +8}\] \[\frac{-5\left(-\pi +8\right)+\left|\pi -8\right|}{-\pi +8}\quad \left(\mathrm{Decimal:\quad }\:-4\right)\]

OpenStudy (decentnabeel):

are you understand @lisa123